OpenStudy (anonymous):

Which of the following constants can be added to x^2 - 3x to form a perfect square trinomial? 1 1/2 2 1/4 4 1/2

OpenStudy (anonymous):

Like what I've told earlier, divide the coefficient of x (in this case, it is -3), then once you get the quotient, square it.

OpenStudy (anonymous):

so divide -3 then square it?

OpenStudy (anonymous):

Oops, I forgot to write that you should divide -3 by 2.

OpenStudy (anonymous):

which is 1.5?

OpenStudy (anonymous):

Keep it in fraction, since your choices are in fraction.

OpenStudy (anonymous):

sqrt 1 1/2

OpenStudy (anonymous):

Why is it there a square root? -3 divided by 2 is -3/2.

OpenStudy (anonymous):

Then square it. \[\left( -\frac{ 3 }{ 2 } \right)^2\]

OpenStudy (anonymous):

divide by 2 then find square root?

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

Not the square root, but you square it. ^2

OpenStudy (anonymous):

which would just be 1 1/2

OpenStudy (anonymous):

\[-\frac{ 3 }{ 2 }\times-\frac{ 3 }{ 2 }=????\]

OpenStudy (anonymous):

What is -3 times -3 (numerator)? What is 2 time 2 (denominator)?

OpenStudy (anonymous):

9/4

OpenStudy (anonymous):

= 2 1/4

OpenStudy (anonymous):

Yes. :)